i need some help can some please help me .And thanks in advance .

How can translate to algebra statement ;do not solve :
Three times the difference of five and a twice a number yields the same result as the same number increased by four .
Thanks for helpimg me .

Found 2 solutions by stanbon, rajagopalan:
Answer by stanbon(29603)

Three times the difference of five and a twice a number yields the same result as the same number increased by four .
-----------
Let the number be "x".
EQUATION:
3(5-2x)=x+4
=======
Cheers,
Stan H.

Answer by rajagopalan(133)

Given : "Three times the difference of five and a twice a number yields the same result as the same number increased by four . "
Solution:
Let the number be x
diff if five = 5-x
3 times of taht = 3(5-x).....Part1
Twice the number =2x ........Part 2
Add part1&Part2 wthch gives = 3(5-x)+2x
Now equate this to number inc by 4 = x+4
The needed eqn is :
3(5-x)+2x =x+4
Okay
You can solve if required.
Good day..